OpenStudy (anonymous):

How can I solve this y=C1e^2x cos3x+C2e^3x sin3x?

OpenStudy (anonymous):

@ikram002p

zepdrix (zepdrix):

\[\Large\rm y=c_1 e^{2x}\cos 3x+c_2 e^{2x}\sin3x\]This is probably the general solution to a second order differential equation. Do you have initial conditions to find a specific solution?

OpenStudy (ikram002p):

solve what lol ? u mean how to find c1, C2 whats the origion equation ?

OpenStudy (anonymous):

zepdrix i don't have the initial conditions :(

zepdrix (zepdrix):

Then this `is` your solution :o It's a general solution. You can't go any further than that without initial conditions D:

OpenStudy (ikram002p):

@ganeshie8

OpenStudy (ikram002p):

yeah or type the origion Qn tht include y'' ,y' ,y

OpenStudy (anonymous):

ahm thankyou for answering my question zepdrix :) I have another question can you please help me to solve ? :)

ganeshie8 (ganeshie8):

I think she wants to eliminate the constants

OpenStudy (anonymous):

ganeshie8 You're right I just want to eliminate the constants

ganeshie8 (ganeshie8):

okay... since you have two constants - you need to dfferentiate it twice

OpenStudy (anonymous):

ganeshie08 can you please help me to solve that question?

ganeshie8 (ganeshie8):

differentiating and solving looks complicated >.< @ikram002p any shortcut ?

OpenStudy (dan815):

you can turn it into a DE to eliminate the parameter constants

OpenStudy (dan815):

like ganeshie already said xD

zepdrix (zepdrix):

That will lead you back to the differential equation that corresponded to this solution. Is that what you're trying to do vanessa?

OpenStudy (anonymous):

y=Ae^2x + Bxe^2x How about this ? Can you solve this?

OpenStudy (anonymous):

zepdrix I just want to eliminate the constants.

OpenStudy (dan815):

a repeated root lead to this solution

OpenStudy (dan815):

you can work back to see what the differential equation must have been

OpenStudy (dan815):

That equation will not have these constants of variation in it yet

OpenStudy (dan815):

what must your characteristic equatoin have been to get these solutions

OpenStudy (anonymous):

This is my first time studying D.E

OpenStudy (dan815):

okay lets look at a homogenous 2nd order linear equation

OpenStudy (anonymous):

dan815 please help me :)

OpenStudy (dan815):

y''+ay'+by=0

OpenStudy (anonymous):

Thanks dan815 :)

OpenStudy (dan815):

we say that y=e^rx is a solution and try to find what r has to be

OpenStudy (dan815):

y'=re^rx y''=r^2e^rx y''+ay'+by=0 e^rx(r^2+ar+b)=0 now this is what your solution was y=Ae^2x + Bxe^2x this happens when you have a reapeated root of r=2

OpenStudy (dan815):

so r^2+ar+b = (r-2)^2

OpenStudy (dan815):

anr (r-2)^2=r^2-4r+4 therefore a= -4 and b=4

OpenStudy (dan815):

your equation must have been y''-4y'+4y=0 This was a possible Differential equation

OpenStudy (dan815):

which no longer has those variation parameters

OpenStudy (anonymous):

I just want to eliminate the constants

OpenStudy (dan815):

we did, this is the only way I can think of without having BVP or IVPs

OpenStudy (anonymous):

Would you please give me a simple way to solve that?

OpenStudy (dan815):

hmm I think they want you to notice that the equation must have been from a repeated root that was the hint

OpenStudy (dan815):

You had to have solved a couple of DEs before to know the general form

OpenStudy (anonymous):

Waaaaaah I don't still understand XD

ganeshie8 (ganeshie8):

may be lets create a table

OpenStudy (dan815):

Your first question complex root, along with its conjugate not repeated* sorry

OpenStudy (anonymous):

that's alright :) but please help me to solve my second question :(

OpenStudy (anonymous):

ganeshie8 my second question is y=Ae^2x+Bxe^2x

ganeshie8 (ganeshie8):

y = Ae^2x + Bxe^2x = e^2x(A + Bx) y' = e^2x(0+B) + 2e^2x(A+Bx) = Be^2x + 2y y'' = 2Be^2x + 2y' -----------------------------

ganeshie8 (ganeshie8):

I have just differentiated the equation twice ^

OpenStudy (dan815):

oh i see :) this way of elimination works too

ganeshie8 (ganeshie8):

we still need to eliminate B from the last equation... think of a clever way to combine those 3 equations... hmmm im still thinking..

OpenStudy (anonymous):

Thanks for your help guys :)

OpenStudy (dan815):

it gets a little long but i can see a way, just start rearranging, b interms of y'' and ' and then the first equation in terms of b and x

ganeshie8 (ganeshie8):

yeah we can subtract last two equations i think

OpenStudy (dan815):

then you will have to find a a linear combination where b and A are 0 or something

ganeshie8 (ganeshie8):

0 : y = Ae^2x + Bxe^2x = e^2x(A + Bx) -2 : y' = e^2x(0+B) + 2e^2x(A+Bx) = Be^2x + 2y 1 : y'' = 2Be^2x + 2y' ----------------------------- - 2y'+y'' = - 4y + 2y'

OpenStudy (dan815):

nice =]

OpenStudy (anonymous):

the answer is -4y+2y' am I right?

ganeshie8 (ganeshie8):

the answer is an equation, don't forget the left hand side !

ganeshie8 (ganeshie8):

- 2y'+y'' = - 4y + 2y' y'' - 4y' + 4y = 0

OpenStudy (anonymous):

ahh ok so the answer should be -2y'+y"=-4y'+2y'

OpenStudy (anonymous):

Ok thanks ganeshie8!!!!!!!!! :)

ganeshie8 (ganeshie8):

yes, and you may make it look neat by putting it in standard form ^

OpenStudy (anonymous):

ganeshie8 and dan815 Thanks for your help guys! Ajah :)

OpenStudy (anonymous):

1 big smile for you guys! :)

OpenStudy (dan815):

yep also note thee are other solutions like any other multiple of the left side

OpenStudy (dan815):

k(y'' - 4y' + 4y) = 0 k E R

OpenStudy (dan815):

since they would be dependant

OpenStudy (dan815):

maybe I shouldnt say other solutons, but different way to say the same solution, because when you are eliminatnig, you might not always end up with the simplest representation

OpenStudy (dan815):

so, your answer on a test might not look the same way, so just rearrange till it does

OpenStudy (anonymous):

dan815 ok so you rearrange again the solution?